Seawater contains 411 ppm calcium (Ca2+). What is the molar concentration (i.e., molarity) of calcium in seawater?
Answer #1

Ans :

ppm stands for parts per million

So 411 ppm calcium = 411 g / 1000 L

Number of moles of calcium = mass / molar mass

= 411 / 40.078 = 10.25 mol

Molarity = mol / volume

= 10.25 / 1000

= 0.0102 M